# CLAP: Isolating Content from Style through Contrastive Learning with Augmented Prompts

### v5

@@ Abstract
Contrastive vision-language models, such as CLIP, have garnered considerable attention for various downstream tasks, mainly due to the remarkable generalization ability of the learned features. However, the features they learn often blend content and style information, which somewhat limits their generalization capabilities under distribution shifts. To address this limitation, we adopt a causal generative perspective for multimodal data and propose contrastive learning with data augmentation to disentangle content features from the original representations. To achieve this, we begin by exploring image augmentation techniques and develop a method to seamlessly integrate them into pre-trained CLIP-like models to extract pure content features. Taking a step further, and recognizing the inherent semantic richness and logical structure of text data, we explore the use of text augmentation to isolate latent content from style features. This enables CLIP-like models’ encoders to concentrate on latent content information, refining the representations learned by pre-trained CLIP-like models. Our extensive experiments across diverse datasets demonstrate significant improvements in zero-shot and few-shot classification tasks, alongside enhanced robustness to various perturbations. These results underscore the effectiveness of our proposed methods in refining vision-language representations and advancing the state of the art in multimodal learning.
Keywords Data Augmentation  $\cdot$ Latent Variables  $\cdot$ Disentanglement
@@ 1 Introduction
Vision-language models, exemplified by CLIP [36], have garnered substantial attention due to their exceptional generalization capabilities, achieved through the learned features obtained by utilizing a cross-modal contrastive loss [36, 20, 25]. However, despite being pre-trained on extensive datasets, CLIP-like models fall short in disentangling latent content information and latent style information. Consequently, they are not immune to spurious correlations, i.e., style-related information is erroneously utilized to predict task-related labels. These limitations become evident in the presence of distribution shifts or adversarial attacks, where spurious correlations often change across different environments. For example: (1) a notable dependence on specific input text prompts has been reported for zero-shot capabilities [48, 47, 21]; (2) performance decline in few-shot scenarios has been observed [36, 13]; and (3) susceptibility to adversarial attacks has been explored [33, 45, 43].
@@
Taking a causal perspective, this work begins with a simple yet effective method—image augmentation—to disentangle content and style information within the learned representations of CLIP-like models. This approach is inspired by recent advancements in theoretical developments in causal representation learning [41], which demonstrate that augmented images can be interpreted as results of soft interventions on latent style variables, as depicted in fig. 1. Such augmentation results in a natural data pair where content information remains unchanged while style information changes. Consequently, using contrastive learning, it becomes feasible to isolate the invariant content information from the variant style information. Motivated by this theoretical advancement, we propose a practical method to incorporate image augmentation into CLIP-like models to extract content information from the originally learned features. Specifically, a disentangled network is designed to fine-tune the pre-trained CLIP model using contrastive loss with image augmentation.
@@
Despite the advancements made in disentangling content and style information from the original features learned by CLIP-like models through image augmentation, we recognize an inherent limitation: it is generally challenging to design adequate image augmentations that ensure all style factors change in an image. Theoretically, disentangling content and style information necessitates changes in all style factors [41]. However, inducing sufficient changes in latent style through image augmentation poses challenges due to the high dimensionality and complexity of style information in image data. Achieving significant style variation via artificially designed image augmentation techniques—such as transforming a photograph of a dog into a sketch while preserving content but dramatically altering style—is notably difficult.
@@
Taking a further step, rather than relying on image augmentation, we explore the use of text augmentation to disentangle latent content and style factors. This shift is motivated by two key observations: (1) vision and language data share the same latent space, and therefore, text augmentation can also be utilized to induce changes in latent style factors; (2) text data inherently possesses high semanticity and logical structure, making it more amenable to property-wise manipulation compared to image data. Consequently, implementing sufficient style changes through text augmentation is more feasible than image augmentation, contributing to isolating content from style information (see fig. 1 for visual comparison). For instance, transforming text from "a photo of a dog" to "a sketch of a dog" is straightforward in the language modality, whereas achieving a similar transformation in image data is challenging. Inspired by these observations, we posit that introducing style variations through text augmentation, as illustrated in fig. 1, provides a more effective approach for learning vision-language content features than relying solely on image augmentation.
In summary, our contributions include:
  1. 1.
    Aimed at disentangling latent content and style factors to refine vision-language features of pre-trained CLIP-like models, we propose contrastive learning with data augmentation to fine-tune the original features from a causal perspective.
  2. 2.
    We present a novel method customized for pre-trained CLIP-like models. This method leverages a disentangled network, which is trained using contrastive learning with image augmentation, to extract latent content features from the learned features provided by the image encoder.
  3. 3.
@@
    We propose Contrastive Learning with Augmented Prompts (CLAP) to extract latent content features from representations of CLIP-like models. It begins by training a disentangled network using the pre-trained text encoder and text augmentation. Subsequently, the trained disentangled network is transferred to the image encoder.
  4. 4.
    Experiments conducted on a large real-world dataset demonstrate the effectiveness of the proposed image and text augmentations in improving zero-shot and few-shot performance, as well as robustness against perturbations.
@@ 2 Related Work
2.1 Contrastive Vision-Language Models
@@
Using a cross-modal contrastive loss, CLIP [36] introduced a scalable contrastive vision-language model by leveraging a large corpus of internet-sourced image-text pairs, demonstrating unprecedented zero-shot learning capabilities and exceptional generalization ability across datasets, while supporting numerous downstream tasks [38]. ALIGN [20] expanded the scale of contrastive vision-language modeling by training on up to one billion image-text pairs and integrating the vision transformer’s self-attention mechanism [11], which further enhanced performance.
Despite their successes, CLIP-like models exhibit sensitivity to input text prompts [48, 21], leading to variable performance across different prompts. Efforts to mitigate this prompt sensitivity through prompt learning and engineering [48, 21, 9, 47, 14] focus on customizing prompts for specific tasks but do not fundamentally enhance CLIP’s representations. Furthermore, CLIP-like models are vulnerable to adversarial attacks [4, 12], with current strategies [33, 45] involving adversarial-natural image pairs to improve resilience.
Our work diverges from these task-specific approaches by aiming to enhance CLIP’s representations from a disentanglement perspective, addressing the aforementioned issues inherent in CLIP-like models.
2.2 Disentangled Representation Learning
Aimed at segregating intrinsic latent factors in data into distinct, controllable representations, disentangled representation learning benefits various applications [44, 40, 24]. Specifically, in classification tasks, it has been shown that improving model performance and robustness against data distribution perturbations can be achieved by more effectively disentangling invariant content variables, without needing to fully identify all intrinsic latent variables [22, 28, 26, 27].
@@ 2.2 Disentangled Representation Learning
Within single modalities, studies such as [49] have illustrated that contrastive learning [7, 16, 18] can potentially reverse the data-generative process, aiding in the separation of representations. Furthermore, [41] suggest that image augmentation can help isolate content variables from the latent space through significant stylistic changes. [19] employs mixture techniques for data augmentation, enabling more abundant cross-modal matches.
Diverging from these methods, our approach focuses on employing text augmentation to disentangle latent content variables, introducing a unique approach to learning refined vision-language representations.
3 A Causal Generative Model for Multimodal Data
To understand pre-trained CLIP-like models, we investigate the underlying causal generative process for vision-language data. We consider the following causal generative model, as depicted in fig. 1. In the proposed model, the shared latent space governing vision and language data is divided into two distinct subspaces: one corresponding to the latent content variables $\mathbf{c}$ and the other to the latent style variables $\mathbf{s}$. The latent content variables are posited to determine the object label $\mathbf{y}$, a relationship corroborated by prior studies [22, 29, 31].
@@ 3 A Causal Generative Model for Multimodal Data
Furthermore, to elucidate the correlation between the latent style variable $\mathbf{s}$ and the object variable $\mathbf{y}$, our model incorporates the premise that the latent content variable $\mathbf{c}$ causally influences the latent style variable $\mathbf{s}$, in concordance with the principles of causal representation learning highlighted in recent literature [10, 29, 41]. Additionally, considering the diversity between image data and text data—where information in image data is typically much more detailed, while information in text data tends to be more logically structured—we posit distinct causal mechanisms for the respective generative processes. Our causal generative model is formulated using structural causal models [2]:
$\mathbf{s}:=\mathbf{g}_{s}(\mathbf{c}),~{}\mathbf{x}:=\mathbf{g}_{x}(\mathbf{c,s}),~{}\mathbf{t}:=\mathbf{g}_{t}(\mathbf{c,s}),~{}\mathbf{y}:=\mathbf{g}_{y}(\mathbf{c}).$ (1)
@@
In Eq. 1, the style variable $\mathbf{s}$ is causally influenced by the content via $\mathbf{g}_{s}$; $\mathbf{x}$ and $\mathbf{t}$ denote visual and textual data, respectively. Both visual and textual data are causally produced by the shared latent variables $\mathbf{c}$ and $\mathbf{s}$ through distinct, reversible generative processes: $\mathbf{g}_{x}$ for images and $\mathbf{g}_{t}$ for text. The label $\mathbf{y}$ of a sample is exclusively determined by the content variable $\mathbf{c}$ via $\mathbf{g}_{y}$. For simplicity, exogenous noise variables are implicitly assumed but not explicitly represented in the formulation of the causal generative model, in line with the common understanding that each latent variable is influenced by exogenous noise.
Recent seminal work in [41] has demonstrated that the latent content variable $\mathbf{c}$ can be identified up to block identifiability (i.e., $\mathbf{c}$ can be isolated from the style variable $\mathbf{s}$) by requiring all latent style variables to change (e.g., soft interventions on all latent style variables). This change can be achieved through image augmentation; that is, the augmented image $\mathbf{\tilde{x}}$ can be interpreted as a generative result of $\mathbf{\tilde{s}}$, which is produced through soft interventions on the original latent style variables $\mathbf{s}$.
@@
Despite this theoretical advancement, the practical implementation of this result within CLIP-like models remains unclear. In this study, we propose a practical method to disentangle content and style information within CLIP-like models by employing image augmentation, as detailed in § 4.1. Moreover, we recognize that implementing sufficient changes in all latent style variables $\mathbf{s}$ through text augmentation is more feasible than through image augmentation. Due to the high semanticity and logical structure of text data, we further explore the use of text augmentation to separate content information from style information, as discussed in § 4.2.
@@ 4 Isolating Content from Style with Data Augmentation
In this section, we propose the use of data augmentation to extract content information from the learned features in pre-trained CLIP-like models. Essentially, data augmentation facilitates the alteration of style factors while preserving content factors. Consequently, leveraging contrastive learning enables the segregation of content information from style information. We explore two distinct forms of data augmentation, namely image augmentation (§ 4.1) and text augmentation (§ 4.2).
@@ 4.1 Isolating Content from Style with Augmented Images
While recent studies (von et al., 2021) have provided theoretical assurance regarding the disentanglement of content and style through contrastive learning with data augmentation, it remains unclear how these findings can be applied to vision-language models. We extend these theoretical results to CLIP-like models as follows.
The theoretical framework suggests using InfoNCE loss [34] to extract content information, as formulated below:
$\mathcal{L}(\mathbf{f};\{\mathbf{x}_{i},\mathbf{\tilde{x}}_{i}\}_{i=1}^{b},\tau)=-\frac{1}{b}\sum\nolimits_{i=1}^{b}\log\frac{\exp\bigl{(}\langle\mathbf{f}(\mathbf{x}_{i}),\mathbf{f}(\mathbf{\tilde{x}}_{i})\rangle/\tau\bigr{)}}{\sum_{j=1}^{b}\exp\bigl{(}\langle\mathbf{f}(\mathbf{x}_{i}),\mathbf{f}(\mathbf{\tilde{x}}_{j})\rangle/\tau\bigr{)}},$ (2)
[…]
@@
Figure 2: Refining CLIP through data augmentation. (a) Training involves a disentangled network $\mathbf{f}_{c}$, utilizing contrastive loss on original and augmented image pairs $\mathbf{x}$ and $\mathbf{\tilde{x}}$, with CLIP’s image encoder $\mathbf{f}^{*}_{x}$ having frozen gradients. (b) More efficient content feature learning is achieved through contrastive learning with augmented text prompts $\mathbf{t}$ and $\mathbf{\tilde{t}}$, using the fixed text encoder $\mathbf{f}^{*}_{t}$ of CLIP. (c) Inference stage: The trained disentangled network $\mathbf{f}^{*}_{c}$ integrates with CLIP’s text and image encoders, $\mathbf{f}^{*}_{t}$ and $\mathbf{f}^{*}_{x}$, to enable zero-shot inference for an input image $\mathbf{x}$ and class names $\mathbf{t}_{1}$ to $\mathbf{t}_{n}$.
@@
We extend this formulation to refine pre-trained vision-language models using contrastive learning with augmented images (hereinafter referred to as "Im.Aug"). As illustrated in fig. 2(a), we train a disentangled network on top of CLIP’s pre-trained image encoder. To improve training efficiency and usability, we freeze the pre-trained image encoder. Based on the InfoNCE loss, the learning objective of Im.Aug is defined as:
@@
$\mathbf{f}^{*}_{c}=\mathop{\mathrm{argmin}}_{\mathbf{f}_{c}}\mathop{\mathbb{E}}_{\{\mathbf{x}_{i}\}_{i=1}^{b}\in\mathcal{D}_{x}}\mathcal{L}(\mathbf{f}_{c}\circ\mathbf{f}^{*}_{x};\{\mathbf{x}_{i},\mathbf{\tilde{x}}_{i}\}_{i=1}^{b},\tau),$ (3)
@@
where $\mathcal{D}_{x}$ denotes the training image dataset, $b$ is the batch size, and $\mathbf{f}_{c}$ is the disentangled network being trained. The pre-trained CLIP image encoder is represented by $\mathbf{f}^{*}_{x}$, with the asterisk "*" indicating that its weights are fixed. The variable $\mathbf{x}_{i}$ refers to an image sampled from $\mathcal{D}_{x}$, and $\mathbf{\tilde{x}}_{i}$ is its augmented view.
@@
The composition of the training dataset $\mathcal{D}_{x}$, the image augmentation techniques used, the structure of the disentangled network $\mathbf{f}_{c}$, and the usage of $\mathbf{f}^{*}_{c}$ after training are detailed in the following subsections.
Data Synthesis and Image Augmentation

### v6

@@ Abstract
Contrastive vision-language models, such as CLIP, have garnered considerable attention for various downstream tasks, mainly due to the remarkable ability of the learned features for generalization. However, the features they learned often blend content and style information, which somewhat limits their generalization capabilities under distribution shifts. To address this limitation, we adopt a causal generative perspective for multimodal data and propose contrastive learning with data augmentation to disentangle content features from the original representations. To achieve this, we begin with exploring image augmentation techniques and develop a method to seamlessly integrate them into pre-trained CLIP-like models to extract pure content features. Taking a step further, recognizing the inherent semantic richness and logical structure of text data, we explore the use of text augmentation to isolate latent content from style features. This enables CLIP-like model’s encoders to concentrate on latent content information, refining the learned representations by pre-trained CLIP-like models. Our extensive experiments across diverse datasets demonstrate significant improvements in zero-shot and few-shot classification tasks, alongside enhanced robustness to various perturbations. These results underscore the effectiveness of our proposed methods in refining vision-language representations and advancing the state-of-the-art in multimodal learning.  [^1] Our code is available at https://github.com/YichaoCai1/CLAP.
Keywords:
Data Augmentation Latent Variables Disentanglement
@@ 1 Introduction
Vision-language models, exemplified by CLIP [36], have garnered substantial attention due to their exceptional generalization capabilities, achieved through the learned features, obtained by utilizing a cross-modal contrastive loss [36, 20, 25]. However, despite being pre-trained on extensive datasets, CLIP-like models fall short in disentangling latent content information and latent style information. Consequently, they are not immune to spurious correlations, i.e., style-related information is erroneously utilized to predict task-related labels. These limitations become evident in the presence of distribution shifts or adversarial attacks where spurious correlations often change across different environments. For examples, (1) a notable dependence on specific input text prompts has been reported for zero-shot capabilities [48, 47, 21]; (2) performance decline in few-shot scenarios has been observed in few-shot learning scenarios [36, 13]; and (3) susceptibility to adversarial attacks has been explored [33, 45, 43].
@@
Taking a causal perspective, this work begin with a simple yet effective method, image augmentation, to disentangle content and style information within the learned representations of CLIP-like models. This approach is inspired by recent advancements in theoretical development in causal representation learning [41], which demonstrate that augmented image can be interpreted as a result of soft interventions on latent style variables, as depicted in Fig. 1. Such augmentation results in a natural data pair where content information remains unchanged while style information changes. Consequently, using contrastive learning, it becomes feasible to isolate the invariant content information from the variant style information. Motivated by this theoretical advancement, we propose a practical method to incorporate image augmentation into CLIP-like models to extract content information from the original learned features. Specifically, a disentangled network is designed to fine-tune the pre-trained CLIP model by using a contrastive loss with image augmentation.
@@
Despite the advancements made in disentangling content and style information from the original features learned by CLIP-like models through image augmentation, we recognize an inherent limitation: it is generally challenging to design adequate image augmentations to ensure all style factors change in an image. Theoretically, disentangling content and style information necessitates changes in all style factors [41]. However, inducing sufficient changes in latent style through image augmentation poses challenges due to the high dimensionality and complexity of style information in image data. Achieving significant style variation via artificially designed image augmentation techniques, such as transforming a photograph of a dog into a sketch while preserving content but dramatically altering style, is notably difficult.
@@
Taking a further step, rather than relying on image augmentation, we explore the use of text augmentation to disentangle latent content and style factors. This shift is motivated by two key observations: 1) Vision and language data share the same latent space. Therefore, text augmentation can also be utilized to induce changes in latent style factors instead of image augmentation. 2) Text data inherently possesses high semanticity and logical structure, making it more amenable to property-wise manipulation compared to image data. Consequently, implementing sufficient style changes through text augmentation is more feasible than image augmentation, contributing to isolating content from style information, see Fig. 1 for visual comparison. For instance, transforming text from "a photo of a dog" to "a sketch of a dog" is straightforward in the language modality, whereas achieving a similar transformation in image data is challenging. Inspired by these observations, we posit that introducing style variations through text augmentation, as illustrated Fig. 1, provides a more effective approach for learning vision-language content features than relying on image augmentation.
@@
In summary, our contributions include: (1) Aimed at disentangling latent content and style factors to refine vision-language features of pre-trained CLIP-like models, we propose constrastive learning with data augmentation to fine tune the original features of pre-trained CLIP-like models from a causal perspective. (2) We present a novel method customized for pre-trained CLIP-like models. This method leverages a disentangled network, which is trained using contrastive learning with image augmentation, to extract latent content features from the learned features provided by image encoder of CLIP-like models. (3) We propose Contrastive Learning with Augmented Prompts (CLAP), to extract latent content features from representations of CLIP-like models. It begins by training a disentangled network using the pre-trained text encoder of CLIP-like models and text augmentation. Subsequently, the trained disentangled network is transferred to the image encoder of CLIP-like models. (4) Experiments conducted on a large real dataset demonstrate the effectiveness of the proposed image augmentation and text augmentation in terms of zero-shot and few-shot performance, as well as robustness against perturbations.
@@ 2 Related Work
2.0.1 Contrastive Vision-Language Models
@@
Using a cross-modal contrastive loss, CLIP [36] revolutionarily introduced a scalable contrastive vision-language model by leveraging a large corpus of internet-sourced image-text pairs, demonstrating unprecedented zero-shot learning capabilities and exceptional generalization ability across datasets and supporting numerous downstream tasks [38]. ALIGN [20] expanded the scale of contrastive vision-language modeling, training on up to one billion image-text pairs while integrating the vision transformer’s self-attention mechanism [11], which further enhanced performance. Despite their successes, CLIP-like models exhibit sensitivity to input text prompts [48, 21], leading to variable performance across different prompts. Efforts to mitigate this prompt sensitivity through prompt learning and engineering [48, 21, 9, 47, 14] focus on customizing prompts for specific tasks but do not fundamentally enhance CLIP’s representations. Furthermore, CLIP-like models are vulnerable to adversarial attacks [4, 12], with current strategies [33, 45] involving adversarial-natural image pairs to improve resilience. Our work diverges from task-specific approaches by aiming to enhance CLIP’s representations from a disentanglement perspective, addressing the aforementioned issues inherent in CLIP-like models.
2.0.2 Disentangled Representation Learning
@@
Aimed at segregating intrinsic latent factors in data into distinct, controllable representations, disentangled representation learning benefits various applications [44, 40, 24]. Specifically, in classification tasks, it’s shown that enhancing the model’s performance and robustness against data distribution perturbations can be achieved by more effectively disentangling invariant content variables, without needing to identify all intrinsic latent variables completely [22, 26, 27, 28]. Within single modalities, studies such as [49] have illustrated that contrastive learning [7, 16, 18] can potentially reverse the data generative process, aiding in the separation of representations. Furthermore, [41] suggest that image augmentation can help isolate content variables from the latent space through significant stylistic changes. [19] employs mixture techniques for data augmentation, enabling more abundant cross-modal matches. Diverging from these methods, our approach focuses on employing text augmentation to disentangle latent content variables, introducing a unique approach to learn refined vision-language representations.
3 A Causal Generative Model for Multi-Modal Data
@@
To understand pretrained CLIP-like models, we investigate the underlying causal generative process for vision-language data. We consider the following causal generative model as depicted in Fig. 1. In the proposed model, the shared latent space ruling vision and language data is divided into two distinct sub-spaces: one corresponding to the latent content variables $\mathbf{c}$ and the other to the latent style variables $\mathbf{s}$. The latent content variables are posited to determine the object label $\mathbf{y}$, a relationship corroborated by prior studies [22, 29, 31]. Furthermore, to elucidate the correlation between the latent style variable $\mathbf{s}$ and the object variable $\mathbf{y}$, our model incorporates the premise that the latent content variable $\mathbf{c}$ causally influences the latent style variable $\mathbf{s}$, in concordance with the principles of causal representation learning highlighted in recent literature [10, 29, 41]. Additionally, considering the diversity between image data and text data, where information in image data is typically much more details while information in text data tends to be more logically structured nature, we posit distinct causal mechanisms for the generation processes. Our causal generative model is formulated as following structural causal models [2]:
$\mathbf{s}:=\mathbf{g_{s}(c)},\quad\mathbf{x}:=\mathbf{g_{x}(c,s)},\quad\mathbf{t}:=\mathbf{g_{t}(c,s)},\quad\mathbf{y}:=\mathbf{g_{y}(c)}.$ (1)
@@
In Eq. 1, the style variable $\mathbf{s}$ is causally influenced by the content via $\mathbf{g_{s}}$; $\mathbf{x}$ and $\mathbf{t}$ denote visual and textual data, respectively. Both visual and textual data are causally produced by the shared latent variables $\mathbf{c}$ and $\mathbf{s}$ through distinct, reversible generative processes: $\mathbf{g_{x}}$ for images and $\mathbf{g_{t}}$ for text data, respectively. The label $\mathbf{y}$ of a sample is exclusively determined by the content variable $\mathbf{c}$ via $\mathbf{g_{y}}$.For simplicity, exogenous noises are implicitly assumed but not explicitly represented in the causal generative model’s formulation, aligning with the common understanding that each latent variable is influenced by exogenous noise.
@@
Recent seminal work in [41] has demonstrated that the latent content variable $\mathbf{c}$ can be identified up to block identifiability (i.e., $\mathbf{c}$ can be isolated from style variable $\mathbf{s}$), by requiring all latent style variables to change (e.g., soft interventions on all latent style variables). This change can be achieved through image augmentation, i.e., the augmented image $\mathbf{\tilde{x}}$ can be interpreted as a generative result of $\mathbf{\tilde{s}}$, which is produced through soft interventions on original latent style variables $\mathbf{s}$. Despite such theoretical advancement, the practical implementation of this theoretical result within CLIP-like models remains unclear. In this study, we propose a practical method to disentangle content and style information within CLIP-like models by employing image augmentation, as detailed in Sec. 4.1. Moreover, we recognize that implementing sufficient changes on all latent style variables $\mathbf{s}$ through text augmentation is more feasible than image augmentation, due to high semanticity and logical structure in text data, we delve into the use of text augmentation to separate content information from style information, as discussed in Sec. 4.2.
@@ 4 Isolating Content from Style with Data Augmentation
In this section, we propose the employment of data augmentation to extract content information from the learned features in pre-trained CLIP-like models. Essentially, data augmentation facilitates the alteration of style factors while preserving content factors. Consequently, leveraging contrastive learning enables the segregation of content information from style information. We delve into two distinct forms of data augmentation, namely image augmentation (Sec. 4.1) and text augmentation (Sec. 4.2).
@@ 4.1 Isolating Content from Style with Augmented Images
While recent studies (von et al., 2021) have offered assurance regarding the disentanglement of content and style through contrastive learning with data augmentation, it remains unclear how these theoretical findings can be applied to the realm of vision-language models. We convert the theoretical findings into CLIP-like models in the following. The theoretical findings suggest using InfoNCE loss [34] to extract content information, as outlined below:
$\mathcal{L}{(\mathbf{f};\{\mathbf{x}_{i},\mathbf{\tilde{x}}_{i}\}_{i=1}^{b},\tau)}=-\frac{1}{b}\sum\nolimits_{i=1}^{b}\mathrm{log}\frac{\mathrm{exp}\left[\langle\mathbf{f}(\mathbf{x}_{i}),\mathbf{f}(\mathbf{\tilde{x}}_{i})\rangle/\tau\right]}{\sum_{j=1}^{b}\mathrm{exp}\left[\langle\mathbf{f}(\mathbf{x}_{i}),\mathbf{f}(\mathbf{\tilde{x}}_{j})\rangle/\tau\right]},$ (2)
where $\{\mathbf{x}_{i}\}_{i=1}^{b}$ represents a batch of $b$ samples from the training dataset, $\mathbf{f}(\mathbf{x}_{i})$ denotes sample $\mathbf{x}_{i}$’s features through model $\mathbf{f}$, $\mathbf{\tilde{x}}_{i}$ is the augmented counterpart of $\mathbf{x}_{i}$, and $\mathrm{\langle\mathbf{z}_{1},\mathbf{z}_{2}\rangle}$ represents the cosine similarity between two feature vectors, $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$, and $\tau$ represents the temperature parameter influencing the loss.
@@
Figure 2: Refining CLIP through data augmentation. (a) Training involves a disentangled network $\mathbf{f_{c}}$, utilizing contrastive loss on original and augmented image pairs $\mathbf{x}$ and $\mathbf{\tilde{x}}$, with CLIP’s image encoder $\mathbf{f^{*}_{x}}$ holding frozen gradients. (b) More efficient content feature learning is achieved through contrastive learning with augmented text prompts $\mathbf{t}$ and $\mathbf{\tilde{t}}$, using the fixed text encoder $\mathbf{f^{*}_{t}}$ of CLIP. (c) Inference stage: The trained disentangled network $\mathbf{f^{*}_{c}}$ integrates with CLIP’s text and image encoders, $\mathbf{f^{*}_{t}}$ and $\mathbf{f^{*}_{x}}$, to enable zero-shot inference for an input image $\mathbf{x}$ and class names $\mathbf{t}_{1}$ to $\mathbf{t}_{n}$.
@@
We extend it to refine pre-trained vision-language models, utilizing contrastive learning with augmented images (hereinafter referred to as "Im.Aug"). As illustrated in Fig. 2(a), we train a disentangled network on top of CLIP’s pre-trained image encoder. To enhance training efficiency and the usability of the proposed method, we freeze the pre-trained image encoder. Based on an InfoNCE loss, the learning objective of Im.Aug is formulated as follows:
@@
$\mathbf{f^{*}_{c}}=\mathop{\mathrm{argmin}}_{\mathbf{f_{c}}}\mathop{\mathbb{E}}_{\{\mathbf{x}_{i}\}_{i=1}^{b}\in\mathcal{D}_{\mathbf{x}}}\mathcal{L}(\mathbf{f_{c}\circ f^{*}_{x}};\{\mathbf{x}_{i},\mathbf{\tilde{x}}_{i}\}_{i=1}^{b},\tau),$ (3)
@@
where $\mathcal{D}_{\mathbf{x}}$ denotes the training image dataset and $b$ represents the batch size, $\mathbf{f_{c}}$ is the disentangled network undergoing training. The pre-trained CLIP image encoder is represented by $\mathbf{f}^{*}_{\mathbf{x}}$, with the asterisk "*" signifying that the model weights remain fixed. The variable $\mathbf{x}_{i}$ refers to an image sampled from $\mathcal{D}_{\mathbf{x}}$, and $\mathbf{\tilde{x}}_{i}$ is its augmented view.
@@
The composition of the training dataset $\mathcal{D}_{\mathbf{x}}$, the image augmentation techniques used, the structure of the disentangled network $\mathbf{f_{c}}$, and the utilization of $\mathbf{f^{*}_{c}}$ post-training are detailed in the following subsections.
4.1.1 Data Synthesis and Image Augmentation